# Scaling Up Liquid-Resistance Liquid-Capacitance Networks for Efficient Sequence Modeling

### v2

@@ 7 Conclusion
In this work, we revisited the potential of nonlinear RNNs in the era of efficient, scalable LSSMs. While LSSMs have seen remarkable success due to their parallelizable structure and computational efficiency, nonlinear RNNs have largely been sidelined due to their inherently sequential nature. However, recent advances, particularly the DEER and ELK methods, have opened the door to parallelizing nonlinear RNNs, thus challenging their long-standing scalability limitation.
Building on these developments, we introduced the liquid-resistance liquid-capacitance nonlinear state-space model (LrcSSM), a novel NSSM architecture that combines the expressive power of bioinspired nonlinear RNNs with the scalability of modern LSSMs. By adapting the ELK method and carefully redesigning the internal structure of LRCs, we enable efficient parallel computation by inherently learning diagonal Jacobian matrices, while still preserving the dynamic richness of nonlinear state updates in biological neurons. Our design allows for exact ELK updates, rather than relying on quasi-approximations.
Our experiments demonstrate that LrcSSM not only matches but often exceeds the performance of leading LSSMs such as LRU, S5, S6, and Mamba, particularly in long-horizon sequence modeling tasks. These results suggest that nonlinear-RNN-based SSMs are not only a feasible solution but can also be competitive, offering a promising direction for future research in sequence modeling.
In summary, this work bridges the gap between the expressive flexibility of nonlinear dynamics and the computational advantages of parallelism, within the LrcSSMs architecture, opening new pathways for scalable, biologically inspired architectures in modern deep learning.
@@ A.1 Stability
where $\rho$ is a user–chosen radius (typically $0.9\!-\!0.99$) enforced by either the tanh-clamp or the negative–softplus–exponential parametrisation.
@@ A.2 Scalability
Let $\text{FLOPs}\;\;\approx\;\;c_{f}\,B\,T\,D\,L$ , be the dominant training cost, where $B$ is the batch size, $T$ the sequence length, $D$ the hidden width, $L$ the network depth, and $c_{f}$ an architecture–specific constant we define (lower for SSMs and higher for Transformers). The single-GPU throughput (tokens s-1 GPU-1) is $\text{throughput}\;\;\approx\;\;\frac{TB}{\text{wall-clock time}}\;$ The memory footprint is the sum of peak activations and model parameters.
Scaling-law [18, 15]. A scaling law is any asymptotic or empirical relation of the form
@@
or a closed-form complexity identity such as $\text{FLOPs}\;\propto\;T\,D.$
[…]
Table 4 summarizes the per–layer cost of the main long-sequence architectures in terms of forward/backward FLOPs, peak activation memory, and parallel depth over the sequence length $T$. Because LrcSSM shares the same $\mathcal{O}(TD)$ compute curve as Mamba but with a smaller constant $c_{f}$ (no low-rank gate, no FFT), we expect it to sit at—or slightly below—the 0.420.420.420.420.450.450.450.45 band. The claim is compatible with existing data: Mamba-3B matches a 6-B Transformer at the same FLOPs [9], and LinOSS shows 2×2$\times$2 × lower NLL than Mamba on $50$ k-token sequences at equal compute [32]. Hence, $\beta\approx 0.42$ is a defensible prior for LrcSSM; a hybrid LrcSSM + local-attention block could plausibly move $\beta$ toward $0.41$.
@@
LrcSSM (ours) $\mathcal{O}(TD)$ $\mathcal{O}(D)$ $\mathcal{O}(\log T)$
@@
For single–GPU throughput $K(T)$, LrcSSM inherits the near-perfect linear behaviour $K(T)\!\propto\!T$ of the scan primitive, with practical speed-ups obtainable through width-$w$ windowing and double buffering that saturate L2 cache bandwidth. Liquid-S4 degrades linearly in latency because it remains sequential, whereas FFT-based S4/Hyena layers incur $\mathcal{O}(T\log T)$ compute and become memory-bound beyond $T\!\approx\!64$k tokens. Hence, for contexts up to $64$k, LrcSSM (and Mamba) are the compute winners; at larger $T$ the FFT models may overtake them in raw FLOPs but pay a significant activation cost.
@@
[…]
Thus, for $T\!\leq\!64\,\mathrm{k}$, LrcSSM and Mamba are compute winners; beyond $64\,\mathrm{k}$, Hyena/S4 win in pure flops but can be memory-bound.
@@ B.2 Hyperparameters
hidden dimension ${16,64,128}$
state-space dimension ${16,64,256}$
number of blocks (#blocks) ${2,4,6}$
@@
Table 6: Hyperparameters used for LrcSSM per dataset.
lr hidden dim. state-space dim. #blocks
Heartbeat $10^{-3}$ 64 64 4
SelfRegulationSCP1 $10^{-3}$ 64 16 2
SelfRegulationSCP2 $10^{-3}$ 128 64 2
EthanolConcentration $10^{-4}$ 128 16 2
MotorImagery $10^{-4}$ 16 16 4
EigenWorms $10^{-4}$ 16 16 4
@@ B.3 Dataset sources
B.4 Additional Remarks on the Datasets
We used the datasets as they were publicly available (i.e., without an additional time dimension). However, this aspect is treated as an additional hyperparameter in the models reported by [32]. We hypothesize that incorporating this dimension could help our model learn even better dependencies, which might be worth investigating in the future.
B.5 Additional Remarks on the Model Design
@@
Appendix C Ablation Studies
@@
We conducted ablation studies to assess the importance of incorporating state-dependency in the state-transition matrix $\mathbf{A}$ and input-transition vector $\mathbf{b}$. Given the extensive hyperparameter search required, we fixed the architecture to 6 layers of SSM blocks, each with 64 states, an input encoding dimension of 64, and a learning rate of $10^{-4}$. As shown in Table 7, the average results indicate that learning both input- and state-dependent transitions yields better performance. We also suggest that future work could treat these dependencies as tunable hyperparameters, as some datasets may benefit from both forms of dependency, while others may perform well with input-dependency alone.
Table 7: Experimentation with different input and state-dependent matrices. Here, we use a fixed configuration with input encoding of 64 and 6 blocks of SSMs with 64 units. We found that excluding state dependency from $\mathbf{A}$ and then from $\mathbf{b}$ too, downgrades performance on average.
@@
Heartbeat 75.0 $\pm$ 2.6 75.0 $\pm$ 1.8 73.0 $\pm$ 2.7
SelfRegulationSCP1 84.8 $\pm$ 2.8 85.0 $\pm$ 2.9 83.1 $\pm$ 1.4
SelfRegulationSCP2 55.4 $\pm$ 7.7 49.6 $\pm$ 5.5 51.4 $\pm$ 2.9
EthanolConcentration 36.1 $\pm$ 1.1 37.6 $\pm$ 3.9 34.2 $\pm$ 2.9
MotorImagery 55.7 $\pm$ 4.1 57.9 $\pm$ 2.9 54.3 $\pm$ 6.0
EigenWorms 85.6 $\pm$ 5.4 85.0 $\pm$ 5.5 86.7 $\pm$ 5.4
Average 65.4 $\pm$ 17.9 65.0 $\pm$ 18.0 63.8 $\pm$ 18.7
Please note that results reported here for LrcSSM, do not match the results of the previous tables because we used a fix setup without hyperparameter tuning, to only focus on the importance of state-dependency and changed the underlying matrix $\mathbf{A}$ and $\mathbf{b}$ of $\dot{x}=A(x,u)x+b(x,u)$. This results in having even better test accuracies reported here for Heartbeat and SelfRegulationSCP2.
@@
We also experimented with complex-valued learnable parameters, focusing on those interacting directly with the state $x$. In particular,we experimented with the parameters $g_{i}^{max,x}$ of $f_{i}^{*}(x_{i},u)$ and $k_{i}^{max,x}$ of $z_{i}^{*}(x_{i},u)$ as defined in Eq.(8) and (9), respectively, as well as their shared sigmoidal channel parameters $a_{i}^{x}$ and $b_{i}^{x}$. These were gradually converted to complex values, and experiments were conducted using a fixed configuration of 6 SSM blocks, each with 64 state dimensions and 64-dimensional encoded input, and a learning rate of $10^{-4}$. As shown in Table 8, we found no significant performance gains on average from using complex-valued parameters. As a result, we opted to use real-valued learnable parameters in our main experiments. Nevertheless, we also evaluated the tuned models with their complex-valued counterparts. The only notable improvement occurred on the MotorImagery dataset, where accuracy increased from 54.3 $\pm$ 3.1 to 58.6 $\pm$ 3.1. Substituting this result into the average accuracy reported in Table 3 would yield 65.6 $\pm$ 18.9, which still ranks our model as the second-best overall.
Table 8: Experimentation with complex valued parameters. Here, we use a fixed configuration with input encoding of 64 and 6 blocks of SSMs with 64 units. We found very similar average performance between real-valued and complex-valued parameters.
@@
(default) with with with
@@
Heartbeat 75.0 $\pm$ 2.6 74.3 $\pm$ 5.2 73.75 $\pm$ 3.2 73.0 $\pm$ 4.0
SelfRegulationSCP1 84.8 $\pm$ 2.8 82.9 $\pm$ 2.7 83.1 $\pm$ 4.2 84.8 $\pm$ 2.2
SelfRegulationSCP2 55.4 $\pm$ 7.7 50.4 $\pm$ 4.4 53.6 $\pm$ 3.6 58.6 $\pm$ 3.5
EthanolConcentration 36.1 $\pm$ 1.1 41.8 $\pm$ 2.1 40.0 $\pm$ 4.5 42.1 $\pm$ 3.6
MotorImagery 55.7 $\pm$ 4.1 53.2 $\pm$ 2.6 53.9 $\pm$ 3.5 52.5 $\pm$ 4.3
EigenWorms 85.6 $\pm$ 5.4 85.0 $\pm$ 6.5 88.3 $\pm$ 5.7 86.1 $\pm$ 6.3
Average 65.4 $\pm$ 17.9 64.6 $\pm$ 16.8 65.4 $\pm$ 17.4 66.2 $\pm$ 16.4

### v6

@@ 7 Conclusion
In this work, we revisited the potential of non-linear RNNs in the era of efficient, scalable linear SSMs. While linear SSMs have seen remarkable success due to their parallelizable structure and computational efficiency, non-linear RNNs have largely been sidelined due to their inherently sequential nature. However, recent advances, particularly the DEER [27] and ELK methods, and their quasi-variants [8] have opened the door to parallelizing non-linear RNNs, thus challenging their long-standing scalability limitation.
Building on these developments, we introduced the liquid-resistance liquid-capacitance non-linear state-space model (LrcSSM), a novel SSM architecture that combines the expressive power of bio-inspired non-linear RNNs with the scalability of modern SSMs. By adapting the parallelization methods and carefully redesigning the internal structure of LRCs, we enable efficient parallel computation by inherently learning diagonal Jacobian matrices, while still preserving the dynamic richness of non-linear state updates in biological neurons. Our design allows for exact parallel updates, rather than relying on quasi-approximations.
Our experiments demonstrate that LrcSSM not only matches but often exceeds the performance of leading linear SSMs such as LRU, S5, S6, and Mamba, particularly in long-horizon sequence modeling tasks. These results suggest that non-linear-RNN-based SSMs are not only a feasible solution but can also be competitive, offering a promising direction for future research in sequence modeling.
In summary, this work bridges the gap between the expressive flexibility of non-linear dynamics and the computational advantages of parallelism, within the LrcSSMs architecture, opening new pathways for scalable, biologically inspired architectures in modern deep learning.
Acknowledgements
M.F. has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034277. Experiments were performed on the dataLab cluster at TU Wien, whose support and infrastructure contributed significantly to these research results. We also thank Liquid AI for providing extra computational resources. Finally, we thank the anonymous NeurIPS reviewers for their valuable feedback and comments, which helped to improve our work.
@@ A.1 Stability
where $\rho$ is a user–chosen radius (typically 0.90.990.9\!-\!0.99) enforced by either the tanh-clamp or the negative–softplus–exponential parametrisation.
@@ A.2 Scalability
Let $\text{FLOPs}\;\;\approx\;\;c_{f}\,B\,T\,D\,L$ , be the dominant training cost, where $B$ is the batch size, $T$ the sequence length, $D$ the hidden width, $L$ the network depth, and $c_{f}$ an architecture–specific constant we define (lower for SSMs and higher for Transformers). The single‐GPU throughput (tokens s-1 GPU-1) is $\text{throughput}\;\;\approx\;\;\frac{TB}{\text{wall‐clock time}}\;$ The memory footprint is the sum of peak activations and model parameters.
Scaling‐law [20, 17]. A scaling law is any asymptotic or empirical relation of the form
@@
or a closed‐form complexity identity such as $\text{FLOPs}\;\propto\;T\,D.$
Recent large‐scale studies like [35] show that $\beta$ depends on the operator’s per-token cost: Dense attention: $\beta\approx 0.48$$0.50$ [20]. Linear-time RNN/SSM (Mamba, Hyena): $\beta\approx 0.42$$0.45$ in $70$ M–77 B runs [10, 34]. Hybrid (recurrence + sparse attention): $\beta$ can reach $0.41$ (MAD pipeline) [35].
Table 3 summarizes the per–layer cost of the main long-sequence architectures in terms of forward/backward FLOPs, peak activation memory, and parallel depth over the sequence length $T$. Because LrcSSM shares the same $\mathcal{O}(TD)$ compute curve as Mamba but with a smaller constant $c_{f}$ (no low-rank gate, no FFT), we expect it to sit at—or slightly below—the 0.420.420.450.45 band. The claim is compatible with existing data: Mamba-3B matches a 6-B Transformer at the same FLOPs [10], and LinOSS shows $2\times$ lower NLL than Mamba on $50$ k-token sequences at equal compute [37]. Hence, $\beta\approx 0.42$ is a defensible prior for LrcSSM; a hybrid LrcSSM + local-attention block could plausibly move $\beta$ toward $0.41$.
@@
LrcSSM (ours) $\mathcal{O}(TD)$ $\mathcal{O}(TD)$ $\mathcal{O}(\log T)$
@@
For single–GPU throughput $K(T)$, LrcSSM inherits the near-perfect linear behaviour $K(T)\!\propto\!T$ of the scan primitive, with practical speed-ups obtainable through width-$w$ windowing and double buffering that saturate L2 cache bandwidth. Liquid-S4 degrades linearly in latency because it remains sequential, whereas FFT-based S4/Hyena layers incur $\mathcal{O}(T\log T)$ compute and become memory-bound beyond $T\!\approx\!64$k tokens. Hence, for contexts up to 6464k, LrcSSM (and Mamba) are the compute winners; at larger $T$ the FFT models may overtake them in raw FLOPs but pay a significant activation cost.
@@
Let $K(T)$ be the wall‑clock time for a single forward pass of length $T$ on one GPU. LrcSSM: $K(T)\!\approx\!\dfrac{c}{\mathrm{SMs}}\;T$ (linear) but can drop to $\approx\dfrac{c}{\mathrm{SMs}}\;\dfrac{T}{w}$ with a width‑$w$ scan and double‑buffering—near‑perfect L2‑cache reuse, where SM is the number of CUDA Streaming Multiprocessors on the GPU, and $c$ a hardware-and-kernel–dependent constant (e.g., time per token per SM). Mamba [10]: same asymptotic, but the fused CUDA kernel shows $\approx 5\times$ higher throughput than a Transformer on $4\,\mathrm{k}$ tokens; on shorter sequences the constant cost of its scan kernel dominates. S4/Hyena (FFT): $\mathcal{O}(T\!\log T)$; cross‑over with linear methods occurs around $T\!\approx\!8$$16\,\mathrm{k}$ on A100s—FlashFFTConv reduces the constant $4\times$$8\times$ [6]. Liquid‑S4 [15]: remains sequential; throughput degrades linearly without remedy.
Thus, for $T\!\leq\!64\,\mathrm{k}$, LrcSSM and Mamba are compute winners; beyond $64\,\mathrm{k}$, Hyena/S4 win in pure flops but can be memory‑bound.
@@ B.2 Hyperparameters
hidden dimension 16,64,128{16,64,128}
state-space dimension 16,64,256{16,64,256}
number of blocks (#blocks) 2,4,6{2,4,6}
@@
Table 5: Optimized hyperparameters used for LrcSSM per dataset.
lr hidden dim. state-space dim. #blocks include time
Heart $10^{-3}$ 64 64 4 False
SCP1 $10^{-3}$ 64 16 2 False
SCP2 $10^{-3}$ 128 64 2 False
Ethanol $10^{-4}$ 128 16 2 False
Motor $10^{-4}$ 16 16 4 False
Worms $10^{-4}$ 64 16 4 False
@@ B.3 Dataset Sources
B.4 Additional Remarks on the Model Design
@@
B.5 Run Times and Number of Iterations
Table 6: Run time in seconds for the considered models for 1000 training steps. Values for the models other than LrcSSM are taken from [37].
NRDE NCDE Log-NCDE LRU S5 Mamba S6 LinOSS-IMEX LinOSS-IM LrcSSM
Heart 9539 1177 826 8 11 34 4 4 7 23
SCP1 1014 973 635 9 17 7 3 42 38 12
SCP2 1404 1251 583 9 9 32 7 55 22 15
Ethanol 2256 2217 2056 16 9 255 4 48 8 15
Motor 7616 3778 730 51 16 35 34 128 11 31
Worms 5386 24595 1956 94 31 122 68 37 90 33
Figure 2: Iterations needed for convergence for LrcSSM on average per dataset.
In Table 6 we present the average run time of the best models for 1000 training steps for each dataset. The values for the models other than LrcSSM are reported from [37]. Figure 2 shows the number of iterations needed for convergence for LrcSSM on average per dataset. As one can see, we do not lose significant runtime due to the extra iterations that are needed for the states to converge through the Newton-steps.
Appendix C Additional Experiments
We conducted additional experiments on the PPG-DaLiA dataset [36], and report the results (baselines are taken from [37]) in Table 7. We ran the hyperparameter tuning following the same protocol as before. LrcSSM achieves performance comparable to other models, demonstrating its competitiveness on this benchmark.
Table 7: Mean squared error (MSE × $10^{-2}$) for different models on PPG-DaLiA. As before, the performance of the models marked by † is reported from [37]. Results are averaged over 5 seeds.
Model MSE $\times 10^{-2}$
NRDE 9.90 $\pm$ 0.97
NCDE 13.54 $\pm$ 0.69
Log-NCDE 9.56 $\pm$ 0.59
LRU 12.17 $\pm$ 0.49
S5 12.63 $\pm$ 1.25
S6 12.88 $\pm$ 2.05
Mamba 10.65 $\pm$ 2.20
LinOSS-IMEX 7.50 $\pm$ 0.46
LinOSS-IM 6.40 $\pm$ 0.23
LrcSSM 10.89 $\pm$ 0.96
Appendix D Generalized Model Design
Any non-linear RNN can be reformulated by considering its underlying differential equations using an explicit first-order Euler integration scheme with a unit time step. We assume the non-linear RNN has the following general form:
$\mathbf{x}_{t}=\mathbf{q}_{t}(\mathbf{x}_{t-1},\mathbf{u}_{t})\,\mathbf{x}_{t-1}+\mathbf{s}_{t}(\mathbf{x}_{t-1},\mathbf{u}_{t})$,
where $\mathbf{q}_{t}=\sigma({\mathbf{f_{q}}}(\mathbf{x}_{t-1},\mathbf{u}_{t}))$ and $\mathbf{s}_{t}=\sigma({\mathbf{f_{s}}}(\mathbf{x}_{t-1},\mathbf{u}_{t}))$ are non-linear, potentially state- and input-dependent gated ($\sigma$) functions, which we aim to express it as:
$\dot{\mathbf{x}}=\mathbf{A}(\mathbf{x},\mathbf{u})\,\mathbf{x}+\mathbf{b}(\mathbf{x},\mathbf{u})$.
To see this, consider:
$\dot{\mathbf{x}}=(\mathbf{q}(\mathbf{x},\mathbf{u})-\mathbf{1})\,\mathbf{x}+\mathbf{s}(\mathbf{x},\mathbf{u})$,
which yields (using Euler integration with $\Delta t=1.0$) the original equation:
$\mathbf{x}_{t}=\mathbf{x}_{t-1}+\dot{\mathbf{x}}\,\Delta t=\mathbf{x}_{t-1}+(\mathbf{q}_{t}(\mathbf{x}_{t-1},\mathbf{u}_{t})-\mathbf{1})\,\mathbf{x}_{t-1}+\mathbf{s}_{t}(\mathbf{x}_{t-1},\mathbf{u}_{t})=\mathbf{q}_{t}(\mathbf{x}_{t-1},\mathbf{u}_{t})\,\mathbf{x}_{t-1}+\mathbf{s}_{t}(\mathbf{x}_{t-1},\mathbf{u}_{t})$.
Thus, in this general formulation: $\mathbf{A}(\mathbf{x},\mathbf{u})=\mathbf{q}(\mathbf{x},\mathbf{u})-\mathbf{1},\quad\mathbf{b}(\mathbf{x},\mathbf{u})=\mathbf{s}(\mathbf{x},\mathbf{u})$.
To obtain the desired efficient model representation, we define:
$\mathbf{A}(\mathbf{x},\mathbf{u})=\mathrm{diag}\left([\sigma(f_{{q}_{1}}(x_{1},\mathbf{u}))-1,\,\ldots,\,\sigma(f_{{q}_{i}}(x_{i},\mathbf{u}))-1,\,\ldots,\,\sigma(f_{{q}_{m}}(x_{m},\mathbf{u}))-1]\right)$
$\mathbf{b}(\mathbf{x},\mathbf{u})=[\,\sigma(f_{{s}_{1}}(x_{1},\mathbf{u})),\,\ldots,\,\sigma(f_{{s}_{i}}(x_{i},\mathbf{u})),\,\ldots,\,\sigma(f_{{s}_{m}}(x_{m},\mathbf{u}))\,]$.
D.1 Example of GRU
The GRU equations are:
$\mathbf{z}_{t}=\sigma\left(\mathbf{W}_{z}\left[\,\mathbf{x}_{t-1},\,\mathbf{u}_{t}\,\right]+\mathbf{b}_{z}\right)=\sigma(\mathbf{f}_{z})$
$\mathbf{r}_{t}=\sigma\left(\mathbf{W}_{r}\left[\,\mathbf{x}_{t-1},\,\mathbf{u}_{t}\,\right]+\mathbf{b}_{r}\right)=\sigma(\mathbf{f}_{r})$
$\mathbf{c}_{t}=\tanh\left(\mathbf{W}_{h}\left[\,\mathbf{r}_{t}\mathbf{x}_{t-1},\,\mathbf{u}_{t}\,\right]+\mathbf{b}_{h}\right)=\tau(\mathbf{f}_{p})$
$\mathbf{x}_{t}=(1-\mathbf{z}_{t})\,\mathbf{x}_{t-1}+\mathbf{z}_{t}\,\mathbf{c}_{t}$
Rewriting the update using the differential form:
$\dot{\mathbf{x}}=\mathbf{z}(-\mathbf{x}+\mathbf{c})=-\mathbf{z}\,\mathbf{x}+\mathbf{z}\,\mathbf{c}$, to check, we get back to the original form (using Euler integration with $\Delta t=1.0$):
$\mathbf{x}_{t}=\mathbf{x}_{t-1}+\dot{\mathbf{x}}\,\Delta t=\mathbf{x}_{t-1}-\mathbf{z}_{t}\mathbf{x}_{t-1}+\mathbf{z}_{t}\mathbf{c}_{t}=(\mathbf{1}-\mathbf{z}_{t})\,\mathbf{x}_{t-1}+\mathbf{z}_{t}\,\mathbf{c}_{t}$.
To express this in the form $\dot{\mathbf{x}}=\mathbf{A}(\mathbf{x},\mathbf{u})\,\mathbf{x}+\mathbf{b}(\mathbf{x},\mathbf{u})$, we set: $\dot{\mathbf{x}}=\mathbf{A}(\mathbf{x},\mathbf{u})\,\mathbf{x}+\mathbf{b}(\mathbf{x},\mathbf{u})=-\mathbf{z}\,\mathbf{x}+\mathbf{z}\,\mathbf{c}$, thus $\mathbf{A}(\mathbf{x},\mathbf{u})=-\mathbf{z}$, and $\mathbf{b}(\mathbf{x},\mathbf{u})=\mathbf{z}\,\mathbf{c}$.
Using our proposed formulation for constructing $\mathbf{A}(\mathbf{x},\mathbf{u})$ and $\mathbf{b}(\mathbf{x},\mathbf{u})$:
$\mathbf{A}(\mathbf{x},\mathbf{u})=\mathrm{diag}\left([\,-\sigma(f_{{z}_{1}}({x}_{1},\mathbf{u})),\,\ldots,\,-\sigma(f_{{z}_{i}}({x}_{i},\mathbf{u})),\,\ldots,\,-\sigma(f_{{z}_{m}}({x}_{m},\mathbf{u}))\,]\right)$,
$\mathbf{b}(\mathbf{x},\mathbf{u})=[\,\sigma(f_{{z}_{1}}({x}_{1},\mathbf{u}))\,\tau(f_{{p}_{1}}({x}_{1},\mathbf{u})),\ \ldots,\ \sigma(f_{{z}_{i}}^{*}({x}_{i},\mathbf{u}))\,\tau(f_{{p}_{i}}({x}_{i},\mathbf{u})),\ \ldots,\ \sigma(f_{{z}_{m}}({x}_{m},\mathbf{u}))\,\tau(f_{{p}_{m}}({x}_{m},\mathbf{u}))\,]$.
This reformulation results in a diagonal Jacobian matrix, enabling more efficient and exact (rather than quasi) computations for the parallelization method.
Appendix E Ablation Studies
For the ablation studies below, due to the extensive hyperparameter search required, we fixed the architecture to 6 layers of SSM blocks, each with 64 states, an input encoding dimension of 64, and a learning rate of $10^{-4}$.
Motivation for LRC and Model with Constant Capacitance.
We considered models of non-spiking neurons, such as those based on electrical and chemical synapses. Electrical synapses can be modeled using continuous-time recurrent neural networks (CT-RNNs), but they tend to show degraded performance compared to models based on chemical synapses. In such models, part of the transition matrix is fixed, as in many state-space models (SSMs) before, and lacks input or state dependency, which offers nothing novel in this regard.
Instead, we focused on models of chemical synapses, the so-called liquid neural networks, where the term liquid refers to the input- and state-dependent dynamics of the transition matrix. These dynamics are non-linear. Within this family, Liquid Time-Constant Networks (LTCs) do not have saturation (or gating) terms in their dynamics of Eq. (4), which is desirable for achieving stable behavior. Fortunately, their saturated (gated) counterpart, Saturated LTCs (STCs), does include such terms as in Eq. (5).
We modified the model of STCs the same way as LRCs, achieving a diagonal Jacobian matrix (without the elastance term indicated in orange in the equations) and ended up with StcSSMs. This model has constant capacitance, as assumed in models derived from chemical synapse dynamics.
We compared this model against LrcSSM across the datasets, using a fixed setup. We found that LrcSSM outperforms StcSSM, highlighting the need for the non-linear membrane capacitance as shown in Table 8.
Table 8: Comparison of StcSSM with LrcSSM (proposed model) across datasets.
StcSSM LrcSSM
Heart 75.7 $\pm$ 5.0 75.0 $\pm$ 2.6
SCP1 78.8 $\pm$ 4.3 84.8 $\pm$ 2.8
SCP2 50.4 $\pm$ 6.9 55.4 $\pm$ 7.7
Ethanol 36.8 $\pm$ 4.0 36.1 $\pm$ 1.1
Motor 53.9 $\pm$ 4.4 55.7 $\pm$ 4.1
Worms 85.0 $\pm$ 4.8 85.6 $\pm$ 5.4
Average 63.4 $\pm$ 7.8 65.4 $\pm$ 8.0
Impact of Model Simplification.
We evaluated the LrcSSM model with a dense Jacobian matrix as well, where $\mathbf{A}$ and $\mathbf{b}$ have all input- and state-dependencies in the model, to compare performance and assess any potential trade-offs. We used the same architectural design as before to ensure a fair comparison.
Table 9: Performance comparison of LrcSSM with the full model resulting in a dense vs. our proposed model with inherently diagonal Jacobian matrix.
LrcSSM-full LrcSSM
dense Jacobian diagonal Jacobian
Heart 72.3 $\pm$ 4.2 75.0 $\pm$ 2.6
SCP1 83.6 $\pm$ 1.9 84.8 $\pm$ 2.8
SCP2 49.6 $\pm$ 5.0 55.4 $\pm$ 7.7
Ethanol 33.9 $\pm$ 1.5 36.1 $\pm$ 1.1
Motor 55.7 $\pm$ 5.7 55.7 $\pm$ 4.1
Worms 84.4 $\pm$ 3.8 85.6 $\pm$ 5.4
Average 63.3 $\pm$ 8.3 65.4 $\pm$ 8.0
These results in Table 9 show that empirically, we do not lose the expressive capabilities of LRCs by constraining the Jacobian matrix to be diagonal. We hypothesize that this is because the strict structure encourages the model to learn more efficient representations, while also enabling exact (rather than approximate) computations due to the design.
@@ Appendix E Ablation Studies
We conducted ablation studies to assess the importance of incorporating state-dependency in the state-transition matrix $\mathbf{A}$ and input-transition vector $\mathbf{b}$. As shown in Table 10, the average results indicate that learning both input- and state-dependent transitions yields better performance. We also suggest that future work could treat these dependencies as tunable hyperparameters, as some datasets may benefit from both forms of dependency, while others may perform well with input-dependency alone.
Table 10: Experimentation with different input and state-dependent matrices. Here, we use a fixed configuration with input encoding of 64 and 6 blocks of SSMs with 64 units. For each dataset, the mean and standard deviation are reported, and the last row presents the mean and standard error aggregated across datasets. We found that excluding state dependency from $\mathbf{A}$ and then from $\mathbf{b}$ too, downgrades performance on average.
@@
Heart 75.0 $\pm$ 2.6 75.0 $\pm$ 1.8 73.0 $\pm$ 2.7
SCP1 84.8 $\pm$ 2.8 85.0 $\pm$ 2.9 83.1 $\pm$ 1.4
SCP2 55.4 $\pm$ 7.7 49.6 $\pm$ 5.5 51.4 $\pm$ 2.9
Ethanol 36.1 $\pm$ 1.1 37.6 $\pm$ 3.9 34.2 $\pm$ 2.9
Motor 55.7 $\pm$ 4.1 57.9 $\pm$ 2.9 54.3 $\pm$ 6.0
Worms 85.6 $\pm$ 5.4 85.0 $\pm$ 5.5 86.7 $\pm$ 5.4
Average 65.4 $\pm$ 8.0 65.0 $\pm$ 8.0 63.8 $\pm$ 8.4
Please note that results reported here for LrcSSM, do not match the results of the previous tables because we used a fix setup without hyperparameter tuning, to only focus on the importance of state-dependency and changed the underlying matrix $\mathbf{A}$ and $\mathbf{b}$ of $\mathbf{\dot{x}}=\mathbf{A}(\mathbf{x},\mathbf{u})\mathbf{x}+\mathbf{b}(\mathbf{x},\mathbf{u})$. This results in having even better test accuracies reported here for Heartbeat and SelfRegulationSCP2.
@@
We also experimented with complex-valued learnable parameters, focusing on those interacting directly with the state $\mathbf{x}$. In particular,we experimented with the parameters $g_{i}^{max,x}$ of $f_{i}^{*}(x_{i},\mathbf{u})$ and $k_{i}^{max,x}$ of $z_{i}^{*}(x_{i},\mathbf{u})$ as defined in Eq.(8) and (9), respectively, as well as their shared sigmoidal channel parameters $a_{i}^{x}$ and $b_{i}^{x}$. These were gradually converted to complex values, and experiments were conducted using a fixed configuration of 6 SSM blocks, each with 64 state dimensions and 64-dimensional encoded input, and a learning rate of $10^{-4}$. As shown in Table 11, we found no significant performance gains on average from using complex-valued parameters. As a result, we opted to use real-valued learnable parameters in our main experiments. Nevertheless, we also evaluated the tuned models with their complex-valued counterparts. The only notable improvement occurred on the MotorImagery dataset, where accuracy increased from 54.3 $\pm$ 3.1 to 58.6 $\pm$ 3.1. Substituting this result into the average accuracy reported in Table 1 would yield 65.6%, which still ranks our model as the second-best overall.
Table 11: Experimentation with complex valued parameters. Here, we use a fixed configuration with input encoding of 64 and 6 blocks of SSMs with 64 units. For each dataset, the mean and standard deviation are reported, and the last row presents the mean and standard error aggregated across datasets. We found very similar average performance between real-valued and complex-valued parameters.
@@
(ours) with with with
@@
Heart 75.0 $\pm$ 2.6 74.3 $\pm$ 5.2 73.75 $\pm$ 3.2 73.0 $\pm$ 4.0
SCP1 84.8 $\pm$ 2.8 82.9 $\pm$ 2.7 83.1 $\pm$ 4.2 84.8 $\pm$ 2.2
SCP2 55.4 $\pm$ 7.7 50.4 $\pm$ 4.4 53.6 $\pm$ 3.6 58.6 $\pm$ 3.5
Ethanol 36.1 $\pm$ 1.1 41.8 $\pm$ 2.1 40.0 $\pm$ 4.5 42.1 $\pm$ 3.6
Motor 55.7 $\pm$ 4.1 53.2 $\pm$ 2.6 53.9 $\pm$ 3.5 52.5 $\pm$ 4.3
Worms 85.6 $\pm$ 5.4 85.0 $\pm$ 6.5 88.3 $\pm$ 5.7 86.1 $\pm$ 6.3
Average 65.4 $\pm$ 8.0 64.6 $\pm$ 7.5 65.4 $\pm$ 7.8 66.2 $\pm$ 7.3